# Efficient preparation of the AKLT State with Measurement-based Imaginary Time Evolution

### v2

@@ Abstract
Quantum state preparation plays a crucial role in several areas of quantum information science, in applications such as quantum simulation, quantum metrology and quantum computing. However, typically state preparation requires resources that scale exponentially with the problem size, due to their probabilistic nature or otherwise, making studying such models challenging. In this article, we propose a method to prepare the ground state of the Affleck-Lieb-Kennedy-Tasaki (AKLT) model deterministically using an measurement-based imaginary time evolution (MITE) approach. By taking advantage of the special properties of the AKLT state, we show that it can be prepared efficiently using the MITE approach. Estimates based on the convergence of a sequence of local projections, as well as direct evolution of the MITE algorithm suggest a constant scaling with respect to the number of AKLT sites, which is an exponential improvement over the naive estimate for conveargence. We show that the procedure is compatible with qubit-based simulators, and show that using a variational quantum algorithm for circuit recompilation, the measurement operator required for MITE can be well approximated by a circuit with a much shallower circuit depth compared with the one obtained using the default Qiskit method.
@@ 1 Introduction
In order to examine the properties of a physical model, the task of state preparation is a crucial step for simulating quantum many-body systems, and has attracted much attention recently [33, 34, 35]. The aim of quantum simulation is to investigate the properties of a particular system, which usually amounts to obtaining the low-lying energy eigenstates of a specified Hamiltonian [36]. Several methods have been proposed for realizing this. First, in an analogue quantum simulation approach [5, 37, 38], the system is physically cooled down, and the Hamiltonian of interest is realized using experimental techniques such as optical lattices and optical tweezers [39, 40, 41]. In a digital quantum simulation approach, the quantum phase estimation [42] algorithm, or other Fourier transform methods [43], can be used to read out of the energy eigenstates of the Hamiltonian. However, in order to study the low energy eigenstates, one must be able to prepare a state with a suitably high overlap with the low energy eigenstates. One approach to prepare this is to use an adiabatic procedure, where one slowly changes the Hamiltonian from a known one to the target Hamiltonian [44]. Another approach that has gained recent popularity are variation approaches, such as the variational quantum eigensolver (VQE) [45, 46] and the quantum approximate optimization algorithm (QAOA) [47]. In these approaches, a measurement-feedback approach is used to obtain a low-energy state of a parameterized circuit. Recently, quantum steering has been used to transform an arbitrary initial state to a final target state [48, 49]. Finally, several methods that mimic imaginary time evolution (ITE) have been developed recently. The variational imaginary time evolution (VITE) [50] method works in a similar way to the measurement-feedback approaches above, and uses a parameterized circuit to find a unitary evolution that best approximates the imaginary time evolution. Another approach is measurement-based imaginary time evolution (MITE), which uses measurements combined with adaptive unitary operations to result in a stochastic, yet deterministic procedure to prepare eigenstates of Hamiltonians [51, 52]. Other recent approaches which deal with non-unitary time evolution either approximate the target non-unitary operators as unitary ones [53], or transform them into unitary operators with additional ancilla qubits [54].
@@
Recently, there has been some interest in preparing the ground state of the Afflect, Kennedy, Lieb and Tasaki (AKLT) model [64, 65]. The AKLT model is a fundamental model in condensed matter physics which has served to understand the physics of fractional excitations at its boundaries [64, 66, 67], as well as the strongly correlated symmetry protected topological (SPT) phase with a Haldane gap [65], due to its exactly solvable nature. Furthermore, the foundational valence-bond solid (VBS) state [68] serves as the ground state for the spin-1 AKLT model with bilinear-biquadratic interactions. This state holds significance as it presented crucial evidence, supporting Haldane’s conjecture [69], which posits that one-dimensional integer-spin Heisenberg models exhibit a gap. Despite its association with one-dimensional systems, the VBS construction can be extended to host higher-dimensional arbitrary lattices [64, 70]. It is also interesting from the point of view of quantum computing since it was shown to be realizable experimentally with photonics systems using cluster states [61, 71, 72], which is the primary resource of measurement-based quantum computing [73, 74, 75, 76]. The AKLT model serves as an excellent testing ground for developing further techniques in quantum simulation on NISQ-era devices. As such, several works have recently discussed creating the state on currently available quantum processors [77, 78, 79], including tensor networks-based approaches [80, 81, 82, 83, 84, 85], and approaches involving dissipation on a qutrit superconducting array [86]. However, those methods either require mid-circuit measurements [78], adiabatic preparation of the state [87], or measurement-based post selection [77, 79] or quantum steering [48]. As a result, it is also very difficult to extract other important quantities from these approaches such as a direct measurement of observables to compute entanglement entropy without any feed-foward processing [78]. Other methods dealing with various non-unitary operators include matrix decompositions [88] and diagonal operators [89]. Postselection makes the methods inefficient for longer chains since the probability of a successful preparation reduces drastically with chain length. Hence, approaches that work without the necessity of postselection or complex operations are in demand, such that they are practically implementable for today’s NISQ-era quantum processors.
@@
In this work, we propose an efficient, deterministic (i.e. no postselection) method to prepare the ground state of the AKLT model. Our approach is based upon the MITE methods introduced in Refs. [51, 52], where the ground state of a given Hamiltonian can be prepared deterministically, using measurements and adaptive unitary operations. In addition to the fundamental interest of the AKLT state, the model is interesting to consider in the context of the MITE algorithm due to the very high preparation efficiency that can be attained. For a general Hamiltonian the ground state preparation for MITE requires an exponential time overhead with system size [51]. Remarkably, in the case of AKLT ground state, this can be reduced to constant time with respect to the chain length $N$. This exponential speedup is thanks to the specific properties of the AKLT model that makes this possible, despite the large Hilbert space and the complexity of the quantum many-body wavefunction. We also show that our approach can be implemented not only on qudit-based devices directly, but also on qubit-based devices. The key procedure of MITE, which involves a measurement in the energy eigenbasis of a given Hamiltonian, can be implemented efficiently using a variational optimization algorithm for circuit recompilation.
@@ 2 The AKLT model
$S_{j}^{x}=\frac{1}{\sqrt{2}}\begin{pmatrix}0&1&0\\ 1&0&1\\ 0&1&0\end{pmatrix},S_{j}^{y}=\frac{i}{\sqrt{2}}\begin{pmatrix}0&-1&0\\ 1&0&-1\\ 0&1&0\end{pmatrix},$ (9)
@@
one achieves the ground state of the Hamiltonian from Eq. (1). In Eq. (16), the states are defined as the eigenstates of the Pauli-$z$ operator $\sigma_{j,l}^{z}|n\rangle_{j,l}=(-1)^{n}|n\rangle_{j,l}$ for $n\in\{0,1\}$. The projective (i.e. measurement) nature of Eq. (17) suggests that it might be compatible with MITE, but in fact the random nature of quantum measurements, and the specific initial state from Eq. (16) that is required makes a reliable preparation more difficult. We explain an approach that is more robust and efficient in the following section.
@@ 3.1 Overall scheme
Due to the fact that the Hamiltonian is the sum of all local projectors on the $S=2$ subspace, the property from Eq. (15) can therefore be applied to the Hamiltonian, and we observe that we also have
@@
Note that this is a specific relation for the AKLT state since $\hat{\mathcal{P}}_{j,j+1}^{S=2}$ do not all commute. Our approach to MITE will therefore be to perform a MITE subroutine on each pair of sites $j$ and $j+1$ locally, and sweep over all pair of sites for many iterations until the fidelity is converged to one with the target state. The purpose of the sweeps is to converge towards a consistent state such as to satisfy Eq. (15) for all $j$. We illustrate the sequence for the preparation of the AKLT state in Fig. 1(b). The MITE subroutine is simultaneously applied to each pair of sites with the same $j$ parity. For example, for the $N=6$ case the $(j,j+1)=(1,2),(3,4),(5,6)$ are applied at the same time. This is followed by $(j,j+1)=(2,3),(4,5),(6,1)$. This parallelization is possible because these operations all commute, as they do not share common spins.
@@ 3.2 MITE subroutine on a pair of sites
where $q\in\{0,1\}$ labels the measurement outcomes. In Ref. [51] it is described how to realize a measurement of this form using an ancilla qubit prepared in the state $|+\rangle_{a}=(|0\rangle_{a}+|1\rangle_{a})/\sqrt{2}$. Applying a unitary evolution $\exp(-i\epsilon\hat{\mathcal{P}}^{S=2}_{j,j+1}\otimes\sigma_{a}^{y})$ and measuring the ancilla qubit in the $z$-basis produces the measurement in Eq. (28). Here, $|S,m\rangle_{j,j+1}$ are angular momentum eigenstates with total spin $S$ and projection $m$ and are eigenstates of $\hat{\mathcal{P}}^{S=2}_{j,j+1}$. The eigenstates consists of two coupled $S=1$ particles on sites $j$ and $j+1$. The eigenvalues of the projection operator are given by
@@ 4.1 The AKLT state preparation with spin-$1$ particles
In this section, we show numerical results of the performance for the AKLT state preparation with spin-$1$ particles. This is a direct preparation of the AKLT state in the spin-$1$ representation of Fig. 1. This can be physically implemented using qutrits [92, 93] on various platorms such as trapped ions [94], superconducting quantum processors [95], silicon carbide [96], as well as silicon-photonic integrated circuits [97].
@@
Figure 2: Partial fidelities using the spin-$1$ representation with $N=4$ between (a) odd pair of sites ($F_{\rm 3,4}$ as an example) and (b) even pair of sites ($F_{\rm 1,4}$ as an example). The three curves are three different runs of the MITE subroutine. The data point is recorded for each measurement step $T$ within each MITE subroutine up to $T=100$. For all panels, the interaction time is chosen to be $\epsilon=0.5$. The factor $\eta$ is chosen to be $\eta=4$ such that $E_{\text{th}}=1/8$.
The partial fidelities are shown in Fig. 2. We observe that there is a fast convergence of the partial fidelities to unity with typically $T\sim 20$ measurements, for all chain lengths that we have tried. Both odd and even parity site pairs converge with a similar number of MITE steps. The trajectory for each particular run is different, due to the stochastic nature of quantum measurements, but every run converges to a state with $F_{j,j+1}=1$. We expect that the convergence occurs at similar timescales regardless of the chain length $N$ because the dimension of the MITE subroutine for a pair is constant in each case. Specifically, there are two sites involved with a dimension of $93^{2}=$ 9. Due to the key property of Eq. (15), the convergence of each MITE subroutine is independent of the chain length.
Figure 3: AKLT state preparation using the spin-$1$ representation with $N=4$. (a) the total state fidelity $F_{\rm tot}$ as a function of MITE subroutine round $r$. For each $r$, MITE routines are performed throughout the whole system. (b) The peak value of the amplitude function as a function of time. We perform $100$ trials of simulations, and randomly choose three of them for the illustration in this figure. The data points are recorded within the iteration for each MITE subroutine round $r$ throughout the whole system up to $r=100$. For all panels, we have used $N_{\rm iter}=30$, and the interaction time is chosen to be $\epsilon=0.5$. The factor $\eta$ is chosen to be $\eta=4$ such that $E_{\text{th}}=1/8$.
The fidelities with respect to the AKLT state are shown in Fig. 3(a)(c)(d). Here we plot the total fidelity after $r$ complete rounds of applications of MITE subroutines applied to the whole chain, as shown in Fig. 1(b). We see that the AKLT state can be prepared in typically $r\sim 20$ rounds. Remarkably, there is little dependence of the number of rounds on different chain lengths, despite the exponential difference in Hilbert space dimensions.
@@
In Fig. 3(b), we also plot the peak value computed using Eq. (32) which is recorded at the end each round. The dashed horizontal lines correspond to the energy eigenvalues given in Eq. (29) multiplied by $\epsilon$. We see that the evolutions eventually converge to the target zero energy values after an initially chaotic regime. For an interaction time $\epsilon=0.5$ the sequence shows good convergence but for smaller values we find the convergence is slower. This is as expected since a shorter interaction time corresponds to a weaker measurement, and the time taken to collapse takes longer. In short, our MITE approach proves to be an efficient and deterministic quantum algorithm to prepare the AKLT state, where the number of rounds required for the total procedure remains constant even for a moderately large system size.
Figure 4: Target AKLT state convergence using direct applications from Eq. (39): (a) the total fidelity $F_{\rm tot}$ as a function of direct application rounds $r$ up to $r=15$. The grey dashed line represents $F_{\rm tot}=0.9$. Lighter to darker colors indicate different system size from $N=3$ to $N=9$. (b) the critical value of rounds $r_{c}$ when $F-{\rm tot}=0.9$ as a function of system size $N$. All critical values of rounds $r_{c}$ were obtained via interpolation of the data from panel (a) by numerically solving $F_{\rm tot}(r)=0.9$.

### v3

@@ Abstract
Quantum state preparation plays a crucial role in several areas of quantum information science, in applications such as quantum simulation, quantum metrology and quantum computing. However, typically state preparation requires resources that scale exponentially with the problem size, due to their probabilistic nature or otherwise, making studying such models challenging. In this article, we propose a method to prepare the ground state of the Affleck-Lieb-Kennedy-Tasaki (AKLT) model deterministically using a measurement-based imaginary time evolution (MITE) approach. By taking advantage of the special properties of the AKLT state, we show that it can be prepared efficiently using the MITE approach. Estimates based on the convergence of a sequence of local projections, as well as direct evolution of the MITE algorithm suggest a constant scaling with respect to the number of AKLT sites, which is an exponential improvement over the naive estimate for convergence. We show that the procedure is compatible with qubit-based simulators, and show that using a variational quantum algorithm for circuit recompilation, the measurement operator required for MITE can be well approximated by a circuit with a much shallower circuit depth compared with the one obtained using the default Qiskit method.
@@ 1 Introduction
In order to examine the properties of a physical model, the task of state preparation is a crucial step for simulating quantum many-body systems, and has attracted much attention recently [33, 34, 35]. The aim of quantum simulation is to investigate the properties of a particular system, which usually amounts to obtaining the low-lying energy eigenstates of a specified Hamiltonian [36]. Several methods have been proposed for realizing this. First, in an analogue quantum simulation approach [5, 37, 38], the system is physically cooled down, and the Hamiltonian of interest is realized using experimental techniques such as optical lattices and optical tweezers [39, 40, 41]. In a digital quantum simulation approach, the quantum phase estimation [42] algorithm, or other Fourier transform methods [43], can be used to read out of the energy eigenstates of the Hamiltonian. However, in order to study the low energy eigenstates, one must be able to prepare a state with a suitably high overlap with the low energy eigenstates. One approach to prepare this is to use an adiabatic procedure, where one slowly changes the Hamiltonian from a known one to the target Hamiltonian [44]. Another approach that has gained recent popularity are variation approaches, such as the variational quantum eigensolver (VQE) [45, 46] and the quantum approximate optimization algorithm (QAOA) [47]. In these approaches, a measurement-feedback approach is used to obtain a low-energy state of a parameterized circuit. Recently, quantum steering has been used to transform an arbitrary initial state to a final target state [48, 49]. Finally, several methods that mimic imaginary time evolution (ITE) have been developed recently. The variational imaginary time evolution (VITE) [50] method works in a similar way to the measurement-feedback approaches above, and uses a parameterized circuit to find a unitary evolution that best approximates the imaginary time evolution. Another approach is measurement-based imaginary time evolution (MITE), which uses measurements combined with adaptive unitary operations to result in a stochastic, yet deterministic procedure to prepare eigenstates of Hamiltonians [51, 52]. Other recent approaches which deal with non-unitary time evolution either approximate the target non-unitary operators as unitary ones [53], or transform them into unitary operators with additional ancilla qubits [54, 55].
@@
Recently, there has been some interest in preparing the ground state of the Afflect, Kennedy, Lieb and Tasaki (AKLT) model [65, 66]. The AKLT model is a fundamental model in condensed matter physics which has served to understand the physics of fractional excitations at its boundaries [65, 67, 68], as well as the strongly correlated symmetry protected topological (SPT) phase with a Haldane gap [66], due to its exactly solvable nature. Furthermore, the foundational valence-bond solid (VBS) state [69] serves as the ground state for the spin-1 AKLT model with bilinear-biquadratic interactions. This state holds significance as it presented crucial evidence, supporting Haldane’s conjecture [70], which posits that one-dimensional integer-spin Heisenberg models exhibit a gap. Despite its association with one-dimensional systems, the VBS construction can be extended to host higher-dimensional arbitrary lattices [65, 71]. It is also interesting from the point of view of quantum computing since it was shown to be realizable experimentally with photonics systems using cluster states [62, 72, 73], which is the primary resource of measurement-based quantum computing [74, 75, 76, 77]. The AKLT model also serves as an excellent benchmarking model for testing quantum simulation on NISQ-era devices and further developing associated techniques. In addition, while the one-dimensional AKLT model is exactly solvable and its ground state can be represented as an MPS, from a technical standpoint, it cannot be generated using a constant-depth local unitary circuit. This is because, in the evolution of such circuits, the correlation between any two local operators becomes negligible beyond the boundaries of a causal light cone [78]. As such, several works have recently discussed creating the state on currently available quantum processors [79, 78, 80], including tensor networks-based approaches [81, 82, 83, 84, 85, 86], and approaches involving dissipation on a qutrit superconducting array [87]. However, those methods either require mid-circuit measurements [78], adiabatic preparation of the state [88], or measurement-based post selection [79, 80] or quantum steering [48]. As a result, it is also very difficult to extract other important quantities from these approaches such as a direct measurement of observables to compute entanglement entropy without any feed-foward processing [78]. Other methods dealing with various non-unitary operators include matrix decompositions [89] and diagonal operators [90]. Postselection makes the methods inefficient for longer chains since the probability of a successful preparation reduces drastically with chain length. Hence, compared with the existing preparations, deterministic quantum algorithms with constant depth of quantum circuits that work without the necessity of postselection or complex operations are in demand, such that they are practically implementable for today’s NISQ-era quantum processors.
@@
In this work, we propose an efficient, deterministic (i.e. no postselection) method to prepare the ground state of the AKLT model regardless of the system sizes considered. Our approach is based upon a unique parallelization of the MITE methods introduced in Refs. [51, 52], where the ground state of a given Hamiltonian can be prepared deterministically, using measurements and adaptive unitary operations. While the AKLT state is well-known to be a simultaneous eigenstate of projection operators, they do not commute with each other, meaning that it cannot be represented as a product states of ground states of the local Hamiltonian. In addition to the fundamental interest of the AKLT state, the model is interesting to consider in the context of the MITE algorithm due to the very high preparation efficiency that can be attained. For a general Hamiltonian the ground state preparation for MITE requires an exponential time overhead with system size [51]. Remarkably, in the case of AKLT ground state, this can be reduced to constant time with respect to the chain length $N$. This exponential speedup is thanks to the specific properties of the AKLT model that makes this possible, despite the large Hilbert space and the complexity of the quantum many-body wavefunction. We also show that our approach can be implemented not only on qudit-based devices directly, but also on qubit-based devices. The key procedure of MITE, which involves a measurement in the energy eigenbasis of a given Hamiltonian, can be implemented efficiently using a variational optimization algorithm for circuit recompilation.
@@ 2 The AKLT model
$S_{j}^{x}=\frac{1}{\sqrt{2}}\begin{pmatrix}0&1&0\\ 1&0&1\\ 0&1&0\end{pmatrix},S_{j}^{y}=\frac{i}{\sqrt{2}}\begin{pmatrix}0&-1&0\\ 1&0&-1\\ 0&1&0\end{pmatrix},$
@@
one achieves the ground state of the Hamiltonian from Eq. (1). In Eq. (6), the states are defined as the eigenstates of the Pauli-$z$ operator $\sigma_{j,l}^{z}|n\rangle_{j,l}=(-1)^{n}|n\rangle_{j,l}$ for $n\in\{0,1\}$. The projective (i.e. measurement) nature of Eq. (7) suggests that it might be compatible with MITE, but in fact the random nature of quantum measurements, and the specific initial state from Eq. (6) that is required makes a reliable preparation more difficult. In the following section, we explain an approach that is more robust and efficient.
@@ 3.1 Overall scheme
Due to the fact that the Hamiltonian is the sum of all local projectors on the $S=2$ subspace, the property of Eq. (5) can therefore be applied to the Hamiltonian, and we observe that we also have
@@
Note that this is a specific relation for the AKLT state since $\hat{\mathcal{P}}_{j,j+1}^{S=2}$ do not all commute. Our approach to MITE will therefore be to perform a MITE subroutine on each pair of sites $j$ and $j+1$ locally, and sweep over all pairs of sites for many iterations until the fidelity is converged to one with the target state. The purpose of the sweeps is to converge towards a consistent state such as to satisfy Eq. (5) for all $j$. We illustrate the sequence for the preparation of the AKLT state in Fig. 1(b). The MITE subroutine is simultaneously applied to each pair of sites with the same $j$ parity. For example, for the $N=6$ case the $(j,j+1)=(1,2),(3,4),(5,6)$ are applied at the same time. This is followed by $(j,j+1)=(2,3),(4,5),(6,1)$. This parallelization is possible because these operations all commute, as they do not share common spins.
@@ 3.2 MITE subroutine on a pair of sites
where $q\in\{0,1\}$ labels the measurement outcomes. In Ref. [51] it is described how to realize a measurement of this form using an ancilla qubit prepared in the state $|+\rangle_{a}=(|0\rangle_{a}+|1\rangle_{a})/\sqrt{2}$. Applying a unitary evolution $\exp(-i\epsilon\hat{\mathcal{P}}^{S=2}_{j,j+1}\otimes\sigma_{a}^{y})$ and measuring the ancilla qubit in the $z$-basis produces the measurement in Eq. (12). Here, $|S,m\rangle_{j,j+1}$ are angular momentum eigenstates with total spin $S$ and projection $m$ and are eigenstates of $\hat{\mathcal{P}}^{S=2}_{j,j+1}$. The eigenstates consist of two coupled $S=1$ particles at sites $j$ and $j+1$. The eigenvalues of the projection operator are given by
@@ 4.1 The AKLT state preparation with spin-$1$ particles
In this section, we show numerical results of the performance for the AKLT state preparation with spin-$1$ particles. This is a direct preparation of the AKLT state in the spin-$1$ representation of Fig. 1. This can be physically implemented using qutrits [93, 94] on various platforms such as trapped ions [95], superconducting quantum processors [96], silicon carbide [97], as well as silicon-photonic integrated circuits [98].
@@
Figure 2: Partial fidelities using the spin-$1$ representation between odd pairs of sites (a) $N=4$ ($F_{\rm 3,4}$ as an example), (c) $N=6$ ($F_{\rm 5,6}$ as an example) and between even pairs of sites (b) $N=4$ ($F_{\rm 1,4}$ as an example), (d) $N=6$ ($F_{\rm 1,6}$ as an example). The darker curves are the averaged partial fidelities calculated over $100$ runs. The shaded area indicates the standard deviations. The data points are recorded for each measurement step $T$ within each MITE subroutine up to $T=100$. For all panels, the interaction time is chosen to be $\epsilon=0.5$. The factor $\eta$ is chosen to be $\eta=4$ such that $E_{\text{th}}=1/8$.
The partial fidelities are shown in Fig. 2. Averaging over 100 runs, we observe that there is a convergence of the partial fidelities to unity typically within $T<100$ measurements, for all chain lengths that we have tried. Both odd and even parity site pairs converge with a similar number of MITE steps. The trajectory for each particular run is different, due to the stochastic nature of quantum measurements, but every run converges to a state with $F_{j,j+1}=1$. We expect that the convergence occurs at similar timescales regardless of the chain length $N$ because the dimension of the MITE subroutine for a pair is constant in each case. Specifically, there are two sites involved with a dimension of $93^{2}=$ 9. Due to the key property of Eq. (5), the convergence of each MITE subroutine is independent of the chain length.
Figure 3: AKLT state preparation using the spin-$1$ representation. The total state fidelity $F_{\rm tot}$ as a function of MITE subroutine round $r$ with (a) $N=6$, and (c) $N=4$. For each $r$, MITE routines are performed throughout the whole system. (b) The peak value of the amplitude function as a function of time with $N=6$. The data points are recorded within the iteration for each MITE subroutine round $r$ throughout the whole system up to $r=100$. (d) Comparison of the total state fidelity $F_{\rm tot}$ for $N=6$ (red), $7$ (blue) and $8$ (green). For all panels, we have used $N_{\rm iter}=30$, and the interaction time is chosen to be $\epsilon=0.5$. The darker curves are the averaged quantities calculated over 100 runs. The shaded area indicates the standard deviations. The factor $\eta$ is chosen to be $\eta=4$ such that $E_{\text{th}}=1/8$.
The fidelities with respect to the AKLT state are shown in Fig. 3(a)(c)(d) for various chain lengths. Here we plot the total fidelity after $r$ complete rounds of applications of MITE subroutines applied to the whole chain, as shown in Fig. 1(b). Similarly to the partial fidelities in Fig. 2, we see that the AKLT state is prepared within $r=100$ rounds, based on the average total fidelity calculated over $100$ runs. Remarkably, there is little dependence of the number of rounds on different chain lengths [Fig. 3(d)], despite the exponential difference in Hilbert space dimensions.
@@
In Fig. 3(b), we also plot the average peak value computed using Eq. (16) which is recorded at the end each round. The average peak values are calculated over $100$ runs. The dashed horizontal lines correspond to the zero energy eigenvalues. We see that the time evolutions eventually almost converge to the target zero energy values around $r\approx 80$ (the final peak value is of the magnitude $10^{-3}$), after an initially chaotic regime. For an interaction time $\epsilon=0.5$ the sequence shows good convergence but for smaller values we find the convergence is slower. This is as expected since a shorter interaction time corresponds to a weaker measurement, and the time taken to collapse takes longer. In short, our MITE approach proves to be an efficient and deterministic quantum algorithm to prepare the AKLT state, where the number of rounds required for the total procedure remains constant even for a moderately large system size.
Figure 4: Target AKLT state convergence using direct applications from Eq. (23): (a) the total fidelity $F_{\rm tot}$ as a function of direct application rounds $r$ up to $r=15$. The grey dashed line represents $F_{\rm tot}=0.9$. Lighter to darker colors indicate different system size from $N=3$ to $N=9$. (b) the critical value of rounds $r_{c}$ when $F_{\rm tot}=0.9$ as a function of system size $N$. All critical values of rounds $r_{c}$ were obtained via interpolation of the data from panel (a) by numerically solving $F_{\rm tot}(r)=0.9$.
Finally, in Fig. 5, we illustrate the effect of noise on the AKLT state preparation with the MITE procedure by showcasing two different types of noise: $Z$ dephasing and the $X$ bit flip noise. The noise is modelled by adding a random rotation gate $\prod_{j=1}^{N}\exp{[i\xi_{j}S_{j}^{\nu}]}$ at each MITE subroutine round $r$, where $\xi_{j}$ is a random number drawn from a normal distribution $\exp{[-\xi^{2}/\sigma^{2}]}$ for each site $j$, and $\nu=z(x)$ for dephasing (bit flip) noise. Such a random phase rotation is equivalent to Lindblad dephasing when averaged over many trajectories [99]. We see that for both dephasing and bit flip noise, when $\xi_{j}$ is significantly small ($\sigma^{2}$ is close to zero), the MITE method is still robust and is able to converge to a successful preparation of the target AKLT state within $r<100$ [Fig. 5]. As for the intermediate strength of the order $\sigma^{2}\approx 10^{-2}$, the total fidelity can still reach almost $90\%$. For all strengths of $\sigma^{2}$ considered, the $X$ bit flip noise more strongly suppresses $F_{\rm tot}$ during the early stage of the algorithm [Fig. 5(b)] compared with the $Z$ dephasing effect. The MITE approach tends to work robustly under such noise since it is a measurement-based method. Once it has converged to the correct state, the states tend to be “locked” at this value with fidelities equal to 1. The stochastic iterative approach of MITE means that the particular trajectory that the initial state takes towards the final state is largely irrelevant — there are many paths that MITE can take towards obtaining the final result. Therefore, the presence of small errors along the way do not affect the overall convergence. However, the noise must not be too large such as to remove the convergence altogether. In the case of our simulations this threshold appears to take place at noise levels of $\sigma^{2}\sim 10^{-2}$.
Figure 5: Effect of noise on the AKLT state preparation. (a) $Z$ dephasing noise by adding a rotation $Z$ gate on each site: $\prod_{j=1}^{N}\exp{[i\xi_{j}S_{j}^{z}]}$, and (b) bit flip noise by adding a rotation $X$ gate on each site: $\prod_{j=1}^{N}\exp{[i\xi_{j}S_{j}^{x}]}$. Here, we set $N=4$, and for each site, $\xi_{j}$ is randomly drawn from a normal distribution $\exp{[-\xi^{2}/\sigma^{2}]}$. The noise is added at the beginning of each MITE subroutine round $r$, and the data is averaged over $100$ runs.